3 electrons are held in a line as shown. The middle one is held in place and the outer two are released. How fast are they moving once they are far away? (q = -1.6x10-19C, m = 9.11x10-31kg)10 um 10 um
Answer #1

from conservation of energy

intially they have some potential energy it is balenced by the kinetic energy

kq1q2/r = 1/2mv^2

v = sqrt(2kq1q2/r*m)

v = sqrt(2*9*10^9*1.6^2*10^-38/10^-5*9.1*10^-31)

v = 7.12*10^3 m/s